# Rigorous Modeling of Droplet Impact Dynamics on Micro-Structured Surfaces: Nonlocal Theory and SPH Simulation of Pancake Bouncing

### v1

@@ 3.2 SPH formulations
While well-established SPH discretization schemes exist for local differential operators [47, 48, 49], this work focuses on developing novel formulations for non-local operators. The SPH framework demonstrates particular advantages for non-local operators due to its intrinsic non-local nature through kernel approximations. Specifically, we have
$\left\langle\frac{\mathbf{f}_{\mathrm{att}}}{\rho}\right\rangle_{i}$ $=\left\langle\frac{4\sigma}{\rho^{2}\varepsilon}\int_{\Omega}\rho(\mathbf{y})\nabla_{\mathbf{x}}W^{\textrm{cs}}(\mathbf{x}-\mathbf{y},R_{1})\,\mathrm{d}\mathbf{y}\right\rangle_{i}$
$:=\frac{4\sigma}{\rho_{0}^{2}\varepsilon}\sum_{j=1}^{N_{f}}m_{j}\nabla_{i}W_{ij}^{\textrm{cs},R_{1}},$ (34)
@@
$\left\langle\frac{\mathbf{F}}{\rho}\right\rangle_{i}$ $=\left\langle-\frac{1}{\rho}\int_{\Gamma}p_{\mathrm{rep}}(\mathbf{x},\mathbf{y})\nabla_{\mathbf{x}}W^{\textrm{wq}}(\mathbf{x}-\mathbf{y},R_{2})\,\mathrm{d}\mathbf{y}\right\rangle_{i}$
$:=-\tilde{\alpha}\rho_{0}\bar{v}_{0}^{2}\sum_{k=1}^{N_{s}}\frac{m_{k}}{\rho_{i}\rho_{k}}\left(1-\frac{|\mathbf{x}_{ik}|}{R_{2}}\right)\nabla_{i}W_{ik}^{\textrm{wq},R_{2}},$ (35)
where $\nabla_{i}W_{ij}^{\textrm{cs},R_{1}}:=\nabla_{\mathbf{x}_{i}}W^{\textrm{cs}}(\mathbf{x}_{i}-\mathbf{x}_{j},R_{1})$ and $\nabla_{i}W_{ij}^{\textrm{wq},R_{2}}:=\nabla_{\mathbf{x}_{i}}W^{\textrm{wq}}(\mathbf{x}_{i}-\mathbf{x}_{j},R_{2})$.
@@
$\frac{\mathrm{d}\rho_{i}}{\mathrm{d}t}=$ $\sum_{j=1}^{N_{f}+N_{s}}m_{j}\mathbf{u}_{ij}\cdot\nabla_{i}W_{ij}^{\textrm{wq}},\quad p_{i}=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho_{i}}{\rho_{0}}\right)^{\gamma}-1\right),$ (36)
$\frac{\mathrm{d}\mathbf{u}_{i}}{\mathrm{d}t}=$ $-\sum_{j=1}^{N_{f}}m_{j}\left(\frac{p_{i}}{\rho_{i}^{2}}+\frac{p_{j}}{\rho_{j}^{2}}+\Pi_{i,j}^{\textrm{art}}\right)\nabla_{i}W_{ij}^{\textrm{wq}}+\frac{4\sigma}{\varepsilon\rho_{0}^{2}}\sum_{j=1}^{N_{f}}m_{j}\nabla_{i}W_{ij}^{\textrm{cs},R_{1}}$
$-\sum_{k=1}^{N_{s}}m_{k}\left(\frac{\tilde{\alpha}\rho_{0}\bar{v}_{0}^{2}}{\rho_{k}\rho_{i}}\frac{R_{2}-|\mathbf{x}_{ik}|}{R_{2}}+\Pi_{i,k}^{\textrm{art}}\right)\nabla_{i}W_{ik}^{\textrm{wq},R_{2}}$
$+\sum_{j=1}^{N_{f}}\frac{m_{j}(\mu_{i}+\mu_{j})}{(\rho_{i}+\rho_{j})^{2}/4}\frac{\mathbf{x}_{ij}\cdot\nabla_{i}W_{ij}^{\textrm{wq}}}{|\mathbf{x}_{ij}|^{2}+0.01(h_{ij})^{2}}\mathbf{u}_{ij}+\mathbf{g}_{i},$ (37)
$\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}=$ $(1-\epsilon)\mathbf{u}_{i}+\epsilon\sum_{j=1}^{N_{f}}m_{j}\mathbf{u}_{ij}W_{ij}^{\textrm{wq}}.$ (38)
Here, $\Pi_{a,b}^{\textrm{art}}$ is artificial viscosity term proposed by Monaghan [47] and
$\Pi_{a,b}^{\textrm{art}}$ $=\begin{cases}\frac{-\alpha c\mu_{ab}}{\bar{\rho}_{ab}}&\mathbf{u}_{ab}\cdot\mathbf{x}_{ab}<0;\\ 0&\mathbf{u}_{ab}\cdot\mathbf{x}_{ab}\geq 0;\end{cases},\mu_{ab}=\frac{h\mathbf{u}_{ab}\cdot\mathbf{x}_{ab}}{\mathbf{x}_{ab}^{2}+0.01h^{2}},\bar{\rho}_{ab}=\frac{\rho_{a}+\rho_{b}}{2}.$ (39)
We set $\epsilon=0.5$ [44] and $\alpha=0.015$ to guarantee the stability of the numerical scheme.
@@ 3.3 Time integration
The Predict-Evaluate-Correct (PEC) method is adopted for time integration, offering high computational efficiency and minimal storage demands. For each time step, one can get $\rho_{i}^{n+1}$, $\mathbf{u}_{i}^{n+1}$, $\mathbf{x}_{i}^{n+1},$ $\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\mathbf{u}_{i}^{n+1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\mathbf{x}_{i}^{n+1/2}}{\mathrm{d}t}$ by given $\rho_{i}^{n}$, $\mathbf{u}_{i}^{n}$, $\mathbf{x}_{i}^{n}$, $\frac{\mathrm{d}\rho_{i}^{n-1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\mathbf{u}_{i}^{n-1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\mathbf{u}_{i}^{n-1/2}}{\mathrm{d}t}$. Firstly, we predict the value at time $t=t_{n+\frac{1}{2}}$ by
$\left\{\begin{array}[]{rl}\rho_{i}^{n+1/2}&=\rho_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\rho_{i}^{n-1/2}}{\mathrm{d}t},\\ \mathbf{u}_{i}^{n+1/2}&=\mathbf{u}_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\mathbf{u}_{i}^{n-1/2}}{\mathrm{d}t},\\ \mathbf{x}_{i}^{n+1/2}&=\mathbf{x}_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\mathbf{x}_{i}^{n-1/2}}{\mathrm{d}t},\end{array}\right$. (43)
@@
$\left\{\begin{array}[]{rl}\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\rho_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\mathbf{u}_{i}^{n+1/2},\mathbf{x}_{i}^{n+1/2}),\\ \frac{\mathrm{d}\mathbf{u}_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\mathbf{u}_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\mathbf{u}_{i}^{n+1/2},\mathbf{x}_{i}^{n+1/2}),\\ \frac{\mathrm{d}\mathbf{x}_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\mathbf{u}_{i}^{n+1/2},\mathbf{x}_{i}^{n+1/2}).\end{array}\right$. (47)
@@
$\left\{\begin{array}[]{rl}\rho_{i}^{n+1}&=\rho_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t},\\ \mathbf{u}_{i}^{n+1}&=\mathbf{u}_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\mathbf{u}_{i}^{n+1/2}}{\mathrm{d}t},\\ \mathbf{x}_{i}^{n+1}&=\mathbf{x}_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\mathbf{x}_{i}^{n+1/2}}{\mathrm{d}t}.\end{array}\right$. (51)
@@ 4 Numerical Validation
$We=\frac{\rho v_{0}^{2}r_{0}}{\sigma},$ (52)
where $\rho=1~\mathrm{g/cm}^{3}$ represents the water density and $\sigma=0.0728~\mathrm{N/m}$ denotes the surface tension coefficient. The fluid properties include dynamic viscosity $\mu=10^{-3}~\mathrm{Pa\cdot s}$ ($10^{-6}~\mathrm{g/(mm\cdot ms)}$) and gravitational acceleration $\mathbf{g}=9.81~\mathrm{m/s}^{2}$. Quantitative validation against the experimental data of [8] (see Supplementary Movies 1–3 in their original work, available at: https://www.nature.com/articles/nphys2980#Sec7) confirms the accuracy of our numerical framework. For direct comparison, the corresponding simulations are provided in Supplementary Movies 1–3, showing consistent spatiotemporal patterns with the experiments.
@@ 4.2 Computational Parameters
The SPH simulation utilizes carefully selected parameters to balance numerical accuracy with computational efficiency. We set the initial particle spacing to $\Delta x=0.06~\mathrm{mm}$, providing sufficient resolution to capture detailed droplet dynamics. The thickness of the surface is set as $\varepsilon=0.06~\mathrm{mm}$. For the kernel function implementation, we employ two distinct characteristic length scales: (1) an interaction radius of $R_{1}=0.26~\mathrm{mm}$ for cohesion simulation, and (2) a smoothing length of $h=R_{2}=0.1625~\mathrm{mm}$ for field variable calculations and repulsive force modeling. The numerical speed of sound is set to $c=10(1~\mathrm{m/s}+v_{0})$, and the time step is dynamically adjusted as $\Delta t=0.062R_{1}/c$ to maintain numerical stability. The model incorporates only one adjustable parameter, the dimensionless substrate parameter $\tilde{\alpha}$, which takes fixed values of 57 (horizontal placement) and 38.73 (tilted configuration). Crucially, for horizontal surfaces, $\tilde{\alpha}$ maintains a constant value regardless of Weber number changes, and its single-parameter nature enables easy determination through simple calibration.
@@ 4.3 Droplet Bouncing on Horizontal Microcone Arrays
4.3.1 Case 1: Low-Weber-Number Impact ($We=7.1$)
The experimental setup consists of a water droplet with diameter $r_{0}$ impacting the microcone array at velocity $v_{0}=0.594$ m/s. Snapshots of the numerical simulations are presented in Figure 4 and compared to the experiment. Experimental observations reveal that under low Weber number conditions, droplets exhibit conventional bouncing behavior on micro-cone arrays, characterized by contraction prior to detachment from the substrate. The numerical simulations (Supplementary Movie 1) demonstrate remarkable temporal consistency with physical experiments throughout the entire process. Figure 5 presents the temporal evolution of the droplet’s horizontal diameter and its vertical height relative to the substrate during the bouncing process. We define $t_{\uparrow}$ as the time instant when the droplet vertical height first reaches zero, and $t_{\textrm{contact}}$ as the moment when the droplet height first attains $3h$. A distinct downward impact is observed around t = 10 ms, consistent with experimental observations. This impact, in fact, provides the droplet with sufficient momentum to detach from the substrate.
Figure 4: Comparison of simulated (bottom) and experimental (top; Copyright 2014 Springer Nature) droplet shapes during droplet bouncing at $We=7.1$. Time instances show initial impact at $t=0$ ms, maximal spreading at $t=4.8$ ms, and complete rebound at $t=16.5$ ms.
Figure 5: Diameter and hight of droplet with $\textrm{We}=7.1$
4.3.2 Case 2: Moderate-Weber-Number Impact ($We=14.1$)
At higher impact energy ($v_{0}=0.837$ m/s), the droplet exhibits more complex dynamics. As illustrated in Figure 6, the droplet deviates from conventional bouncing behavior and exhibits a counterintuitive sequence: detachment from the substrate occurs prior to maximum spreading. The close agreement between numerical snapshots and experimental observations demonstrates the remarkable simulation capability of the proposed method. Furthermore, Figure 7 presents the temporal variations of droplet radius and vertical height obtained from numerical simulations (Supplementary Movie 2). A clear sequence is observed: after detaching from the substrate, the droplet continues to spread until reaching its maximum radius before subsequent contraction.
Figure 6: Comparison of simulated (bottom) and experimental (top; Copyright 2014 Springer Nature) droplet shapes during droplet bouncing at $We=14.1$. Time instances show initial impact at $t=0$ ms, complete rebound at $t=3.4$ ms and maximal spreading at $t=4.6$ ms.
Figure 7: Diameter and hight of droplet with $\textrm{We}=14.1$
4.4 Dynamic Evolution Analysis ($We\in[7,24]$)

### v2

@@ 3.2 SPH formulations
While well-established SPH discretization schemes exist for local differential operators [monaghan2005smoothed, liu2003smoothed, Smoothed000274685600002], this work focuses on developing novel formulations for non-local operators. The SPH framework demonstrates particular advantages for non-local operators due to its intrinsic non-local nature through kernel approximations. Specifically, we have
$\left\langle\frac{\boldsymbol{f}_{\mathrm{att}}}{\rho}\right\rangle_{i}$ $=\left\langle\frac{4\sigma}{\rho^{2}\varepsilon}\int_{\Omega}\rho(\boldsymbol{y})\nabla_{\boldsymbol{x}}W^{\text{cs}}(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{att}})\,\mathrm{d}\boldsymbol{y}\right\rangle_{i}$
(32) $:=\frac{4\sigma}{\rho_{0}^{2}\varepsilon}\sum_{j=1}^{N_{f}}m_{j}\nabla_{i}W_{ij}^{\text{cs},R_{\mathrm{att}}},$
@@
$\left\langle\frac{\boldsymbol{F}}{\rho}\right\rangle_{i}$ $=\left\langle-\frac{1}{\rho}k_{0}\int_{\Gamma}\nabla_{\boldsymbol{x}}W^{\text{new}}(\boldsymbol{x}-\boldsymbol{y},R_{\mathrm{rep}})\,\mathrm{d}\boldsymbol{y}\right\rangle_{i}$
(33) $:=-k_{0}\sum_{k=1}^{N_{s}}\frac{m_{k}}{\rho_{i}\rho_{k}}\nabla_{i}W_{ik}^{\text{new},R_{\mathrm{rep}}},$
where $\nabla_{i}W_{ij}^{\text{cs},R_{\mathrm{att}}}:=\nabla_{\boldsymbol{x}_{i}}W^{\text{cs}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j},R_{\mathrm{att}})$ and $\nabla_{i}W_{ij}^{\text{new},R_{\mathrm{rep}}}:=\nabla_{\boldsymbol{x}_{i}}W^{\text{new}}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j},R_{\mathrm{rep}})$.
@@
(34) $\frac{\mathrm{d}\rho_{i}}{\mathrm{d}t}=$ $\sum_{j=1}^{N_{f}+N_{s}}m_{j}\boldsymbol{u}_{ij}\cdot\nabla_{i}W_{ij}^{\text{wq}},\quad p_{i}=\frac{c^{2}\rho_{0}}{\gamma}\left(\left(\frac{\rho_{i}}{\rho_{0}}\right)^{\gamma}-1\right),$
$\frac{\mathrm{d}\boldsymbol{u}_{i}}{\mathrm{d}t}=$ $-\sum_{j=1}^{N_{f}}m_{j}\left(\frac{p_{i}}{\rho_{i}^{2}}+\frac{p_{j}}{\rho_{j}^{2}}+\Pi_{i,j}^{\text{art}}\right)\nabla_{i}W_{ij}^{\text{wq}}+\frac{4\sigma}{\varepsilon\rho_{0}^{2}}\sum_{j=1}^{N_{f}}m_{j}\nabla_{i}W_{ij}^{\text{cs},R_{\mathrm{att}}}$
$-\sum_{k=1}^{N_{s}}m_{k}\left(\frac{k_{0}}{\rho_{k}\rho_{i}}+\Pi_{i,k}^{\text{art}}\right)\nabla_{i}W_{ik}^{\text{new},R_{\mathrm{rep}}}$
(35) $+\sum_{j=1}^{N_{f}}\frac{m_{j}(\mu_{i}+\mu_{j})}{(\rho_{i}+\rho_{j})^{2}/4}\frac{\boldsymbol{x}_{ij}\cdot\nabla_{i}W_{ij}^{\text{wq}}}{|\boldsymbol{x}_{ij}|^{2}+0.01(h_{ij})^{2}}\boldsymbol{u}_{ij}+\boldsymbol{g}_{i},$
(36) $\frac{\mathrm{d}\boldsymbol{x}_{i}}{\mathrm{d}t}=$ $(1-\epsilon)\boldsymbol{u}_{i}+\epsilon\sum_{j=1}^{N_{f}}m_{j}\boldsymbol{u}_{ij}W_{ij}^{\text{wq}}.$
Here, $\Pi_{a,b}^{\text{art}}$ is artificial viscosity term proposed by Monaghan [monaghan2005smoothed] and
(37) $\Pi_{a,b}^{\text{art}}$ $=\begin{cases}\frac{-\alpha c\mu_{ab}}{\bar{\rho}_{ab}}&\boldsymbol{u}_{ab}\cdot\boldsymbol{x}_{ab}<0;\\ 0&\boldsymbol{u}_{ab}\cdot\boldsymbol{x}_{ab}\geq 0;\end{cases},\mu_{ab}=\frac{h\boldsymbol{u}_{ab}\cdot\boldsymbol{x}_{ab}}{\boldsymbol{x}_{ab}^{2}+0.01h^{2}},\bar{\rho}_{ab}=\frac{\rho_{a}+\rho_{b}}{2}.$
We set $\epsilon=0.5$ [monaghan1994simulating] and $\alpha=0.015$ to guarantee the stability of the numerical scheme.
@@ 3.3 Time integration
The Predict-Evaluate-Correct method is adopted for time integration, offering high computational efficiency and minimal storage demands. For each time step, one can get $\rho_{i}^{n+1}$, $\boldsymbol{u}_{i}^{n+1}$, $\boldsymbol{x}_{i}^{n+1},$ $\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\boldsymbol{u}_{i}^{n+1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\boldsymbol{x}_{i}^{n+1/2}}{\mathrm{d}t}$ by given $\rho_{i}^{n}$, $\boldsymbol{u}_{i}^{n}$, $\boldsymbol{x}_{i}^{n}$, $\frac{\mathrm{d}\rho_{i}^{n-1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\boldsymbol{u}_{i}^{n-1/2}}{\mathrm{d}t}$, $\frac{\mathrm{d}\boldsymbol{u}_{i}^{n-1/2}}{\mathrm{d}t}$. Firstly, we predict the value at time $t=t_{n+\frac{1}{2}}$ by
(41) $\left\{\begin{array}[]{rl}\rho_{i}^{n+1/2}&=\rho_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\rho_{i}^{n-1/2}}{\mathrm{d}t},\\ \boldsymbol{u}_{i}^{n+1/2}&=\boldsymbol{u}_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\boldsymbol{u}_{i}^{n-1/2}}{\mathrm{d}t},\\ \boldsymbol{x}_{i}^{n+1/2}&=\boldsymbol{x}_{i}^{n}+\frac{\Delta t}{2}\cdot\frac{\mathrm{d}\boldsymbol{x}_{i}^{n-1/2}}{\mathrm{d}t},\end{array}\right$.
@@
(45) $\left\{\begin{array}[]{rl}\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\rho_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\boldsymbol{u}_{i}^{n+1/2},\boldsymbol{x}_{i}^{n+1/2}),\\ \frac{\mathrm{d}\boldsymbol{u}_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\boldsymbol{u}_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\boldsymbol{u}_{i}^{n+1/2},\boldsymbol{x}_{i}^{n+1/2}),\\ \frac{\mathrm{d}\boldsymbol{x}_{i}^{n+1/2}}{\mathrm{d}t}&=\frac{\mathrm{d}\boldsymbol{x}_{i}}{\mathrm{d}t}(\rho_{i}^{n+1/2},\boldsymbol{u}_{i}^{n+1/2},\boldsymbol{x}_{i}^{n+1/2}).\end{array}\right$.
@@
(49) $\left\{\begin{array}[]{rl}\rho_{i}^{n+1}&=\rho_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\rho_{i}^{n+1/2}}{\mathrm{d}t},\\ \boldsymbol{u}_{i}^{n+1}&=\boldsymbol{u}_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\boldsymbol{u}_{i}^{n+1/2}}{\mathrm{d}t},\\ \boldsymbol{x}_{i}^{n+1}&=\boldsymbol{x}_{i}^{n}+\Delta t\cdot\frac{\mathrm{d}\boldsymbol{x}_{i}^{n+1/2}}{\mathrm{d}t}.\end{array}\right$.
@@ 4 Numerical Validation
(50) $\textit{We}=\frac{\rho v_{0}^{2}r_{0}}{\sigma},$
where $\rho=1~\mathrm{g/cm}^{3}$ represents the water density and $\sigma=0.0728~\mathrm{N/m}$ denotes the surface tension coefficient. The fluid properties include dynamic viscosity $\mu=10^{-3}~\mathrm{Pa\cdot s}$ ($10^{-6}~\mathrm{g/(mm\cdot ms)}$) and gravitational acceleration $\boldsymbol{g}=9.81~\mathrm{m/s}^{2}$. Quantitative validation against the experimental data of [liu2014pancake] (see Supplementary Movies 1–3 in their original work, available at: https://www.nature.com/articles/nphys2980#Sec7) confirms the accuracy of our numerical framework. For direct comparison, the corresponding simulations are provided in Supplementary Movies 1–3, showing consistent spatiotemporal patterns with the experiments.
@@ 4.2 Computational Parameters
The SPH simulation utilizes carefully selected parameters to balance numerical accuracy with computational efficiency. We set the initial particle of fluid spacing to $\Delta x_{f}=0.06~\mathrm{mm}$, providing sufficient resolution to capture detailed droplet dynamics. We set the initial particle of solid spacing to $\Delta x_{s}=0.02~\mathrm{mm}$, providing sufficient resolution to capture the microstructure. The thickness of the surface is set as $\varepsilon=0.06$ mm. For the kernel function implementation, we employ two distinct characteristic length scales: (1) an interaction radius of $R_{\mathrm{att}}=0.26~\mathrm{mm}$ for cohesion simulation, and (2) a smoothing length of $h=R_{\mathrm{rep}}=0.1625~\mathrm{mm}$ for field variable calculations and repulsive force modeling. This configuration ensures the long-range characteristics of attractive interactions while maintaining computational efficiency. The numerical speed of sound is set to $c=10(1~\mathrm{m/s}+v_{0})$, and the time step is dynamically adjusted as $\Delta t=0.062R_{\mathrm{att}}/c$ to maintain numerical stability. The model incorporates only one adjustable parameter, parameter $k_{0}$, which takes fixed values of 0.0178 $\rm g/(mm\cdot ms^{2})$. The key point is that the $k_{0}$ value remains constant regardless of changes in the Weber number, and its single-parameter nature allows it to be easily determined through simple calibration.
@@ 4.3 Droplet Bouncing on Horizontal Microcone Arrays
4.3.1 Case 1: Low-Weber-Number Impact ($\textit{We}=7.1$)
The experimental setup consists of a water droplet with a diameter $r_{0}$ impacting the microcone array at a velocity $v_{0}=0.594$ m/s. Snapshots of the numerical simulations are presented in Figure 4 and are compared to the experiment. Experimental observations reveal that under low Weber number conditions, droplets exhibit conventional bouncing behavior on micro-cone arrays, characterized by contraction prior to detachment from the substrate. The numerical simulations (Supplementary Movie 1) demonstrate remarkable temporal consistency with physical experiments throughout the entire process. Figure 5 (a) presents the temporal evolution of the droplet’s horizontal diameter and its vertical height relative to the substrate during the bouncing process. We define $t_{\uparrow}$ as the time instant when the droplet vertical height first reaches zero, and $t_{\text{contact}}$ as the moment when the droplet height first attains $R_{\mathrm{rep}}$. A distinct downward impact is observed around $t=10$ ms, consistent with experimental observations. This impact, in fact, provides the droplet with sufficient momentum to detach from the substrate.
Figure 4: Comparison of simulated (bottom) and experimental (top; Copyright 2014 Springer Nature) droplet shapes during droplet bouncing at $\textit{We}=7.1$. Time instances show initial impact at $t=0$ ms, maximal spreading at $t=4.8$ ms, and complete rebound at $t=16.5$ ms.
[…]
Figure 5: Diameter and hight of droplet with $\textit{We}=7.1$ and $\textit{We}=14.1$
4.3.2 Case 2: Moderate-Weber-Number Impact ($\textit{We}=14.1$)
At higher impact energy ($v_{0}=0.837$ m/s), the droplet exhibits more complex dynamics. As illustrated in Figure 6, the droplet deviates from conventional bouncing behavior and exhibits a counterintuitive sequence: detachment from the substrate occurs prior to maximum spreading. The close agreement between numerical snapshots and experimental observations demonstrates the remarkable simulation capability of the proposed method. Furthermore, Figure 5 (b) presents the temporal variations of droplet radius and vertical height obtained from numerical simulations (Supplementary Movie 2). A clear sequence is observed: after detaching from the substrate, the droplet continues to spread until reaching its maximum radius before subsequent contraction.
Figure 6: Comparison of simulated (bottom) and experimental (top; Copyright 2014 Springer Nature) droplet shapes during droplet bouncing at $\textit{We}=14.1$. Time instances show initial impact at $t=0$ ms, complete rebound at $t=3.4$ ms and maximal spreading at $t=4.6$ ms.
4.4 Dynamic Evolution Analysis ($\textit{We}\in[7,24]$)